OpenStudy (anonymous):

Solve and check. 6[2 - (2x - 4)] = 2(6 - 3x)

OpenStudy (anonymous):

3(2-(2x-4))=6-3x 6-6x+12=6-3x -6x+12=-3x -3x=-12 x=4

OpenStudy (anonymous):

now the checking part: The left side - using 4 instead of "x": 6(2-(2*4-4)) = 6(2-(4))=-12 The right side - using 4 instead of "x": 2(6-3*4)=2*(6-12)=2*(-6)=-12 both sides are equal -12, so it checks out.